OpenStudy (anonymous):

given : f'(x+1) / f'(x+2) = 1 + ( f(x+1)/f(x+2) ) f(1)=0 and f(2)=1, then ( f(x) / f(x+1) )= ? a)log f(x+1) b)log f(x) c)(e^f(x+1)) - e d)(e^f(x)) - 1

OpenStudy (anonymous):

bookmark

OpenStudy (anonymous):

@mukushla @ganeshie8 @shubhamsrg

OpenStudy (shubhamsrg):

got it maybe.. is it log f(x+1) ??

OpenStudy (anonymous):

yes that correct !! how'd you get that ?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@Callisto @campbell_st @hartnn

OpenStudy (anonymous):

i just got that it cant be (b) as per hit and trial i.e. substituting x= 1 there...but x=1 matches with rest 3 options.. hmm

OpenStudy (anonymous):

Look, this is quite reasonable to solve equation:\[\int\limits \frac{f'(x)}{f(x)} = \ln f(x)\] while \[\int\limits \frac{f'(x+1)}{f(x+1)} = \ln f(x+1)\] etc.

OpenStudy (anonymous):

Now - to get "there" just MULTIPLY BOTH SIDES by f'(x+2) and DIVIDE BOTH SIDES by f(x+1). Then you get to shape of formulas integrable by the above method

OpenStudy (anonymous):

@Aryang ? You asked this , did u not?

hartnn (hartnn):

@Mikael what about right side?? u get f(x+1)/ f'(x+1) for which there is no standard formula. and what about f'(x+2)/f(x+1) ?

OpenStudy (anonymous):

Count your "horses" once again - this is NOT the correct right side AFTER THE OPERATIONS I DESCRIBED

OpenStudy (anonymous):

well,, how do we go about integrating f'(x+2)/f(x+1) ?

OpenStudy (anonymous):

@amistre64 @satellite73

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@shubhamsrg @shubhamsrg @shubhamsrg @shubham.bagrecha

OpenStudy (anonymous):

arey koi to help karo !! i mean please help anybody ?

OpenStudy (anonymous):

atleast you guys bookmark it,,so that you may be in the discussion..

OpenStudy (anonymous):

someone plz post complete solution Mikael or Shubhamsrg

OpenStudy (shubhamsrg):

sorry couldnt come online whole day yesterday,, here's my solution :

OpenStudy (shubhamsrg):

consider this : d/dx ( f(x+1) / f(x+2) ) = [f'(x+1) f(x+2) - f'(x+2) f(x+1) ]/ (f(x+2) )^2 =f'(x+2) [ f'(x+1)/f'(x+2) . f(x+2) - f(x+1) ] / (f(x+2) )^2 =f'(x+2) [( 1 + ( f(x+1)/f(x+2) )) . f(x+2) - f(x+1) ] /(f(x+2))^2 =f'(x+2) [ f(x+2) + f(x+1) -f(x+1) ] / (f(x+2) )^2 d/dx ( f(x+1) / f(x+2)) = f'(x+2)/f(x+2) => d( f(x+1) / f(x+2) )= f'(x+2)/f(x+2) dx integrating both sides, we have ( f(x+1) / f(x+2) ) = ln (f(x+2)) + C for x=0, C=0 =>( f(x+1) / f(x+2) ) = ln (f(x+2)) if we substitute x-1 in place of x, we get ( f(x) / f(x+1) ) = ln(f(x+1)) hope i didnt make any fallacy somewhere!!

OpenStudy (anonymous):

excellent @shubhamsrg !!thank you very much!! that was really genius of you!!